# Computational Hypergraph Discovery, a Gaussian Process framework for connecting the dots

### v1

@@ 4. A Gaussian Process method for Type 3 problems
=$\mathcal{M}=\Bigg{\{}x\in\mathbb{R}^{3}\Bigg{|}\begin{cases}x_{1}+x_{2}+3x_{3}-2&=0\\ x_{1}-x_{2}+x_{3}&=0\end{cases}\Bigg{\}}\,,$ (4.2)
@@
=$\mathcal{M}=\Bigg{\{}x\in\mathbb{R}^{3}\Bigg{|}\begin{cases}x_{1}&=1-2x_{3}\\ x_{2}&=1-x_{3}\end{cases}\Bigg{\}}\,,$ (4.3)
@@
where $\operatorname{Span}(v_{r+1},\ldots,v_{d+1})$ is the zero-eigenspace of $C_{N}$. Here we write $\lambda_{1}\geq\cdots\geq\lambda_{r}>0=\lambda_{r+1}=\cdots=\lambda_{d+1}$ for the eigenvalues of $C_{N}$ (in decreasing order), and $v_{1},\ldots,v_{d+1}$ for the corresponding eigenvectors ($C_{N}v_{i}=\lambda_{i}v_{i}$). The proposed approach extends to the noisy case (when the data points are perturbations of elements of the manifold) by simply replacing the zero-eigenspace of the covariance matrix by the linear span of the eigenvectors associated with eigenvalues that are smaller than some threshold $\epsilon>0$, i.e., by approximating $\mathcal{M}$ with (4.5) where $r$ is such that $\lambda_{1}\geq\cdots\geq\lambda_{r}\geq\epsilon>\lambda_{r+1}\geq\cdots\geq\lambda_{d+1}$. In this affine setting (4.5) allows us to estimate $\mathcal{M}$ directly without RKHS norm minimization/regularization as discussed in Sec. 2 because linear regression does not require regularization in the sufficiently large data regime. Furthermore the process of pruning ancestors (also discussed in Sec. 2) can be replaced by that of identifying sparse elements $v\in\operatorname{Span}(v_{r+1},\ldots,v_{d+1})$ such that $v_{i}=1$.
@@
=$\mathcal{H}=\big{\{}f(x)=v^{T}\psi(x)\mid v\in\mathbb{R}^{d_{\mathcal{S}}}\big{\}}\,.$ (4.6)
@@
=$\mathcal{H}=\Big{\{}f(x)=v_{0}+\sum_{i}v_{i}x_{i}+\sum_{i\leq j}v_{i,j}x_{i}x_{j}\mid v\in\mathbb{R}^{d_{\mathcal{S}}}\Big{\}}\,,$ (4.8)
and, in the large data regime ($N>d_{\mathcal{S}}$), identifying quadratic dependencies between variables becomes equivalent to (1) adding nodes to the hypergraph corresponding to secondary variables obtained from primary variables $x_{i}$ through known functions (for (4.7), these secondary variables are the quadratic monomials $x_{i}x_{j}$, see Fig. 17.(a)), and (2) identifying affine dependencies between the variables of the augmented hypergraph. The problem can, therefore, be reduced to the affine case discussed in Sec. 4.1. Indeed, as in the affine case, the manifold can then be approximated in the regime where the number of data points is larger than the dimension $d_{\mathcal{S}}$ of the feature map by (4.5), where $v_{r},\ldots,v_{N}$ are the eigenvectors of $C_{N}=$(4.4) whose eigenvalues are zero (noiseless case) or smaller than some threshold $\epsilon>0$ (noisy case).
Figure 17. Feature map generalization
Furthermore, the hypergraph representation of the manifold is equivalent to a feature map generalization of Row Echelon Form Reduction to nonlinear systems of equations. For instance, choosing $x_{3}$ as the dependent variable and $x_{1},x_{2}$ as the free variables, =$\mathcal{M}=\{x\in\mathbb{R}^{3}\mid x_{3}-5x_{1}^{2}+x_{2}^{2}-x_{1}x_{2}=0\}$ can be represented as in Fig. 17.(b) where the round node represents the concatenated variable $(x_{1},x_{2})$ and the red arrow represents a quadratic function. The generalization also enables the representation of implicit equations by selecting secondary variables as free variables. For instance, selecting $x_{3}^{2}$ as the free variable and $x_{1},x_{2}$ as the free variables, =$\mathcal{M}=\{x\in\mathbb{R}^{3}\mid x_{1}^{2}+x_{2}^{2}+x_{3}^{2}-1=0\}$ can be represented as in Fig. 17.(c).
@@
This feature-map extension of the affine case discussed in Sec. 4.2 can evidently be generalized to arbitrary degree polynomials and to other basis functions. However, as the dimension $d_{\mathcal{S}}$ of the range of the feature map $\psi$ increases beyond the number $N$ of data points, the problem becomes underdetermined: the data only provides partial information about the manifold, i.e., it is not sufficient to uniquely determine the manifold. Furthermore, if the dimension of the feature map is infinite, then we are always in that low data regime, and we have the additional difficulty that we cannot directly compute with that feature map. On the other hand, if $d_{\mathcal{S}}$ is finite (i.e., if the dictionary of basis functions is finite), then some elements of $\mathcal{F}$ (some constraints defining the manifold $\mathcal{M}$) may not be representable or well approximated as equations of the form $v^{T}\psi(x)=0$. To address these conflicting requirements, we need to kernelize and regularize the proposed approach (as done in interpolation).
@@
for the kernel defined by $\psi$ and observe that $(\mathcal{H},\big{\langle}\cdot,\cdot\big{\rangle}_{\mathcal{H}})$ is the RKHS defined by the kernel $K$ (which is assumed to contain $\mathcal{F}$ in Problem 1). Observe in particular that for $f=v^{T}\psi\in\mathcal{H}$, $K$ satisfies the reproducing property
@@
We will now show that the feature-map PCA variant of Sec. 4.1 (characterizing the subspace of $f\in\mathcal{H}$ such that $f(X)=0$) can be kernelized as a variant of kernel PCA [23]. To describe this write $K(X,X)$ for the $N\times N$ matrix with entries $K(X_{i},X_{j})$. Write $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{r}>0$ for the nonzero eigenvalues of $K(X,X)$ indexed in decreasing order and write $\alpha_{\cdot,i}$ for the corresponding unit-normalized eigenvectors, i.e.
@@
When the feature map $\psi$ is infinite-dimensional, the data only provides partial information about the constraints defining the manifold in the sense that $f(X)=0$ or equivalently $f(\phi)=0$ is a necessary but not sufficient condition for the zero level set of $f$ to be a valid constraint for the manifold (for $f$ to be such that $f(x)=0$ for all $x\in\mathcal{M}$). So we are faced with the following problems: (1) How to regularize? (2) How do we identify free and dependent variables? (3) How do we identify valid constraints for the manifold? The proposed solution will be based on the Kernel Mode Decomposition (KMD) framework introduced in [30].
@@
Now, let us return to our original manifold approximation problem 1 in the kernelized setting of Sec. 4.3.1. Given the data $X$ we cannot regress an element $f\in\mathcal{F}$ directly since the minimizer of $\|f\|_{K}^{2}+\gamma^{-1}\|f(X)\|_{\mathbb{R}^{N}}^{2}$ is the null function. To identify the functions $f\in\mathcal{F}$, we need to decompose them into modes that can be interpreted as a generalization of the notion of free and dependent variables. To describe this, assume that the kernel $K$ can be decomposed as the additive kernel
@@
Then $\mathcal{H}_{K}=\mathcal{H}_{K_{a}}+\mathcal{H}_{K_{s}}+\mathcal{H}_{K_{z}}$ implies that for all function $f\in\mathcal{H}_{K}$, $f$ can be decomposed as $f=f_{a}+f_{s}+f_{z}$ with $(f_{a},f_{s},f_{z})\in\mathcal{H}_{a}\times\mathcal{H}_{s}\times\mathcal{H}_{z}$.
Example 4.4.
@@
that is the sum of a linear kernel, a quadratic kernel and a fully nonlinear kernel. Take $K_{a}$ to be the part of the linear kernel that depends only on $x_{1}$, i.e.,
@@
  • Given a function $g_{a}$ in the RKHS $\mathcal{H}_{K_{a}}$ defined by the kernel $K_{a}$ is there a function $f_{s}$ in the the RKHS $\mathcal{H}_{K_{s}}$ defined by the kernel $K_{s}$ such that $g_{a}(x)\approx f_{s}(x)$ for $x\in\mathcal{M}$?
  • Given a function $g_{a}\in\mathcal{H}_{K_{a}}$ is there a function $f$ in the RKHS $\mathcal{H}_{K}$ defined by the kernel $K$ such that $f(x)\approx 0$ for $x\in\mathcal{M}$ and such that its $f_{a}$ mode is $-g_{a}$ and its $f_{z}$ mode is zero?
Then, the natural answer to the questions and to those of Sec. 4.3.3 is to identify the modes of the constraint $f=f_{a}+f_{s}+f_{z}\in\mathcal{H}$ (such that $f(x)\approx 0$ for $x\in\mathcal{M}$ ) such that $f_{a}=-g_{a}$ and $f_{z}=0$ by selecting $f_{s}$ to be the minimizer of the following variational problem
@@
Taking $g_{a}(x)=x_{1}$ for our running example 4.4, the questions of Sec. 4.3.3 are, as illustrated in Sec. 2, equivalent to asking whether there exists a function $f_{s}\in\mathcal{H}_{K_{s}}$ that does not depend on $x_{1}$ (since $K_{s}$ does not depend on $x_{1}$) such that
@@
As our running example 4.4, we can decompose $K_{s}=$(4.33) as the sum of an affine kernel, a quadratic kernel, and a fully nonlinear kernel, i.e., $m=3$, $K_{s,1}=1+\beta_{1}\sum_{i\not=1}x_{i}x_{i}^{\prime}$, $K_{s,2}=\beta_{2}\sum_{i\leq j,i,j\not=1}x_{i}x_{j}x_{i}^{\prime}x_{j}^{\prime}$ and $K_{s,3}=\beta_{3}\prod_{i\not=1}(1+k(x_{i},x_{i}^{\prime}))$.
@@
Then, we can order these sub-modes from most active to least active and create a new kernel $K_{s}$ by removing the least active modes from the signal and adding them to the mode that is set to be zero. To describe this, let $\pi(1),\cdots,\pi(m)$ be an ordering of the modes by their activation, i.e., $\|f_{s,\pi(1)}\|_{K_{s,\pi(1)}}^{2}\geq\|f_{s,\pi(2)}\|_{K_{s,\pi(2)}}^{2}\geq\cdots$.
@@
For our running example 4.4, where we try to identify the ancestors of the variable $x_{1}$, if the sub-mode associated with the variable $x_{2}$ is found to be least active, then we can try to remove $x_{2}$ from the list of ancestors and try to identify $x_{1}$ as a function of $x_{3}$ to $x_{d}$. This is equivalent to selecting $K_{a}(x,x^{\prime})=\beta_{1}x_{1}x_{1}^{\prime}$,
@@
The Sec. 4.3.4 and 4.3.5 approach to determining the list of ancestors of a given node is to use a fixed threshold (e.g., $\tau=0.5$) to prune nodes. In this approach, we propose an alternative approach that mimics the strategy employed in Principal Component Analysis (PCA) for deciding which modes should be kept and which ones should be removed. The PCA approach is to order the modes in decreasing order of eigenvalues/variance and (1) either keep the smallest number modes holding/explaining a given fraction (e.g., $90\%$) of the variance in the data, (2) or use an inflection point/sharp drop in the decay of the eigenvalues to select which modes should be kept. Here, we propose a similar strategy. First we employ an alternative determination of the least active mode: we iteratively remove the mode that leads to the smallest increase in noise-to-signal ratio, i.e., we remove the mode $t$ such that,
@@
For our running example 4.4 in which we try to find the ancestors of the variable $x_{1}$ this is equivalent to removing the variables or node $t$ whose removal leads to the smallest loss in signal-to-noise ratio (or increase in noise-to-signal ratio) by selecting
@@
Next, we iterate this process, and we plot (a) the noise-to-signal ratio, and (b) the increase in noise-to-signal ratio as a function of the number of ancestors ordered according to this iteration. Fig. 12 illustrates this process and shows that the removal of an essential node leads to a sharp spike in increase in the noise-to-signal ratio (the noise-to-signal ratio jumps from approximately 50-60% to 99%). The identification of this inflection point can be used as a method for effectively and reliably pruning ancestors (see Sec. 2.6 for a detailed illustration).
@@
2:Learned hypergraph
3:$D$leftarrow NormalizeData($D$) // Normalize the data
4:for node $\in V$ do
[…]
6:         SetOfAncestors $\leftarrow$ All other nodes
@@
8:         if SignalToNoiseRatio $>\tau$ then choose that kernel and exit the for loop
@@
12:     while SignalToNoiseRatio $>\tau$ do // Prune ancestors
@@
15:         if SignalToNoiseRatio $>\tau$ then Remove that ancestor
@@
Our overall method is summarized in the pseudocode Alg. 1 and Alg. 2 that we will now describe. Alg. 1 takes the data $D$ (encoded into the matrix $X$ in Sec. 4.3 and 6) and the set of nodes $V$ as an input and produces, as described in Sec. 4.3, for each node $i\in V$ its set of minimal ancestors $A_{i}$ and the simplest possible function $f_{i}$ such that $x_{i}\approx f_{i}\big{(}(x_{j})_{j\in A_{i}}\big{)}$. It employs the default threshold of $0.5$ on the signal-to-noise ratios for its operations. Line 3 normalizes the data (via an affine transformation) so that the samples $X_{i}$ are of mean zero and variance $1$. Given a node with index $i=1$ in Line 4 ($i$ runs through the set of nodes, and we select $i=1$ for ease of presentation), the command in Line 5 refers to selecting a signal kernel of the form $K_{s}=$(4.33) (where $k$ is selected to be a vanilla RBF kernel such as Gaussian or Matérn), with $1\geq\beta_{1}>0=\beta_{2}=\beta_{3}$ for the linear kernel, $1\geq\beta_{1}\geq\beta_{2}>0=\beta_{3}$ for the quadratic kernel and $01\geq\beta_{1}\geq\beta_{2}\geq\beta_{3}>$ 0 for the fully nonlinear (interpolative) kernel. The ComputeSignalToNoiseRatio function in Line 7 computes the signal-to-noise ratio with $g_{a}(x)=x_{1}$ and with the kernel selected in Line 5. The value of $\gamma$ is selected automatically by maximizing the variance of the histogram of eigenvalues of $D_{\gamma}$ as described in Sec. 6.2 (with the kernel $K=K_{s}=$(4.33) selected in Line 5 and $Y=g_{a}(X)$ with $g_{a}(x)=x_{1}$). The value of $\gamma$ is re-computed whenever a node is removed from the list of ancestors, and $K_{s}$ is nonlinear. Lines 13, 14 and 15 are described in Sec. 4.3.5. They correspond to iteratively identifying the ancestor node $t$ contributing the least to the signal and removing that node from the set of ancestors of the node $1$ if the removal of that node $t$ does not send the signal-to-noise ratio below the default threshold $0.5$.
@@
2:Learned hypergraph
3:$D$leftarrow NormalizeData($D$) // Normalize the data
4:for node $\in V$ do
[…]
@@
8:         if SignalToNoiseRatio $>\tau$ then choose that kernel and exit the for loop
@@
13:     SetOfAncestors$(q)$leftarrow All other nodes
@@
17:         SetOfAncestors$(q-1)$leftarrow SetOfAncestors($q$) $\setminus$ LeastImportantAncestor
18:         $q$leftarrow $q-1$
@@
20:     $q^{\dagger}$leftarrow Inflection point in ($q\rightarrow$ NoiseToSignalRatio$(q)$) or spike in ($q\rightarrow$ NoiseToSignalRatio$(q)$ NoiseToSignalRatio$(q-1)$)
21:     FinalSetOfAncestors $\leftarrow$ SetOfAncestors$(q^{\dagger})$

### v2

@@ 5. A Gaussian Process method for Type 3 problems
$\mathcal{M}=\Bigg\{x\in\mathbb{R}^{3}\Bigg|\begin{cases}x_{1}+x_{2}+3x_{3}-2&=0\\ x_{1}-x_{2}+x_{3}&=0\end{cases}\Bigg\}\,$, (5.2)
@@
$\mathcal{M}=\Bigg\{x\in\mathbb{R}^{3}\Bigg|\begin{cases}x_{1}&=1-2x_{3}\\ x_{2}&=1-x_{3}\end{cases}\Bigg\}\,$, (5.3)
@@
where $\operatorname{Span}(v_{r+1},\ldots,v_{d+1})$ is the zero-eigenspace of $C_{N}$. Here we write $\lambda_{1}\geq\cdots\geq\lambda_{r}>0=\lambda_{r+1}=\cdots=\lambda_{d+1}$ for the eigenvalues of $C_{N}$ (in decreasing order), and $v_{1},\ldots,v_{d+1}$ for the corresponding eigenvectors ($C_{N}v_{i}=\lambda_{i}v_{i}$). The proposed approach extends to the noisy case (when the data points are perturbations of elements of the manifold) by simply replacing the zero-eigenspace of the covariance matrix by the linear span of the eigenvectors associated with eigenvalues that are smaller than some threshold $\epsilon>0$, i.e., by approximating $\mathcal{M}$ with (5.5) where $r$ is such that $\lambda_{1}\geq\cdots\geq\lambda_{r}\geq\epsilon>\lambda_{r+1}\geq\cdots\geq\lambda_{d+1}$. In this affine setting (5.5) allows us to estimate $\mathcal{M}$ directly without RKHS norm minimization/regularization because linear regression does not require regularization in the sufficiently large data regime. Furthermore the process of pruning ancestors can be replaced by that of identifying sparse elements $v\in\operatorname{Span}(v_{r+1},\ldots,v_{d+1})$ such that $v_{i}=1$.
Figure 12. Feature map generalization
@@
$\mathcal{H}=\big\{f(x)=v^{T}\psi(x)\mid v\in\mathbb{R}^{d_{\mathcal{S}}}\big\}\,.$ (5.6)
@@
$\mathcal{H}=\Big\{f(x)=v_{0}+\sum_{i}v_{i}x_{i}+\sum_{i\leq j}v_{i,j}x_{i}x_{j}\mid v\in\mathbb{R}^{d_{\mathcal{S}}}\Big\}\,$, (5.8)
and, in the large data regime ($N>d_{\mathcal{S}}$), identifying quadratic dependencies between variables becomes equivalent to (1) adding nodes to the hypergraph corresponding to secondary variables obtained from primary variables $x_{i}$ through known functions (for (5.7), these secondary variables are the quadratic monomials $x_{i}x_{j}$, see Fig. 12.(a)), and (2) identifying affine dependencies between the variables of the augmented hypergraph. The problem can, therefore, be reduced to the previous affine case. Indeed, as in the affine case, the manifold can then be approximated in the regime where the number of data points is larger than the dimension $d_{\mathcal{S}}$ of the feature map by (5.5), where $v_{r},\ldots,v_{N}$ are the eigenvectors of $C_{N}=$(5.4) whose eigenvalues are zero (noiseless case) or smaller than some threshold $\epsilon>0$ (noisy case).
Furthermore, the hypergraph representation of the manifold is equivalent to a feature map generalization of Row Echelon Form Reduction to nonlinear systems of equations. For instance, choosing $x_{3}$ as the dependent variable and $x_{1},x_{2}$ as the free variables, $\mathcal{M}=\{x\in\mathbb{R}^{3}\mid x_{3}-5x_{1}^{2}+x_{2}^{2}-x_{1}x_{2}=0\}$ can be represented as in Fig. 12.(b) where the round node represents the concatenated variable $(x_{1},x_{2})$ and the red arrow represents a quadratic function. The generalization also enables the representation of implicit equations by selecting secondary variables as free variables. For instance, selecting $x_{3}^{2}$ as the free variable and $x_{1},x_{2}$ as the free variables, $\mathcal{M}=\{x\in\mathbb{R}^{3}\mid x_{1}^{2}+x_{2}^{2}+x_{3}^{2}-1=0\}$ can be represented as in Fig. 12.(c).
@@
This feature-map extension of the previously discussed affine case can evidently be generalized to arbitrary degree polynomials and to other basis functions. However, as the dimension $d_{\mathcal{S}}$ of the range of the feature map $\psi$ increases beyond the number $N$ of data points, the problem becomes underdetermined: the data only provides partial information about the manifold, i.e., it is not sufficient to uniquely determine the manifold. Furthermore, if the dimension of the feature map is infinite, then we are always in that low data regime, and we have the additional difficulty that we cannot directly compute with that feature map. On the other hand, if $d_{\mathcal{S}}$ is finite (i.e., if the dictionary of basis functions is finite), then some elements of $\mathcal{F}$ (some constraints defining the manifold $\mathcal{M}$) may not be representable or well approximated as equations of the form $v^{T}\psi(x)=0$. To address these conflicting requirements, we need to kernelize and regularize the proposed approach (as done in interpolation).
@@
for the kernel defined by $\psi$ and observe that (,,)($\mathcal{H}$$\big\langle\cdot,\cdot\big\rangle_{\mathcal{H}}$) is the RKHS defined by the kernel $K$ (which is assumed to contain $\mathcal{F}$ in Problem 1). Observe in particular that for $f=v^{T}\psi\in\mathcal{H}$, $K$ satisfies the reproducing property
@@
We will now show that the previous feature-map PCA variant (characterizing the subspace of $f\in\mathcal{H}$ such that $f(X)=0$) can be kernelized as a variant of kernel PCA [24]. To describe this write $K(X,X)$ for the $N\times N$ matrix with entries $K(X_{i},X_{j})$. Write $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{r}>0$ for the nonzero eigenvalues of $K(X,X)$ indexed in decreasing order and write $\alpha_{\cdot,i}$ for the corresponding unit-normalized eigenvectors, i.e.
@@
When the feature map $\psi$ is infinite-dimensional, the data only provides partial information about the constraints defining the manifold in the sense that $f(X)=0$ or equivalently $f(\phi)=0$ is a necessary but not sufficient condition for the zero level set of $f$ to be a valid constraint for the manifold (for $f$ to be such that $f(x)=0$ for all $x\in\mathcal{M}$). So we are faced with the following problems: (1) How to regularize? (2) How do we identify free and dependent variables? (3) How do we identify valid constraints for the manifold? The proposed solution will be based on the Kernel Mode Decomposition (KMD) framework introduced in [31] (which shares conceptual foundations with Smoothing Spline ANOVA [47]).
@@
Now, let us return to our original manifold approximation problem 1 in the kernelized setting of (5.11). Given the data $X$ we cannot regress an element $f\in\mathcal{F}$ directly since the minimizer of $\|f\|_{K}^{2}+\gamma^{-1}\|f(X)\|_{\mathbb{R}^{N}}^{2}$ is the null function. To identify the functions $f\in\mathcal{F}$, we need to decompose them into modes that can be interpreted as a generalization of the notion of free and dependent variables. To describe this, assume that the kernel $K$ can be decomposed as the additive kernel
@@
Then $\mathcal{H}_{K}=\mathcal{H}_{K_{a}}+\mathcal{H}_{K_{s}}+\mathcal{H}_{K_{z}}$ implies that for all function $\in\mathcal{H}_{K}, ff$ can be decomposed as $f=f_{a}+f_{s}+f_{z}$ with $(f_{a},f_{s},f_{z})\in\mathcal{H}_{a}\times\mathcal{H}_{s}\times\mathcal{H}_{z}$.
Example 1.
@@
that is the sum of a linear kernel, a quadratic kernel, and a fully nonlinear kernel. Take $K_{a}$ to be the part of the linear kernel that depends only on $x_{1}$, i.e.,
@@
  • Given a function $g_{a}$ in the RKHS $\mathcal{H}_{K_{a}}$ defined by the kernel $K_{a}$ is there a function $f_{s}$ in the RKHS $\mathcal{H}_{K_{s}}$ defined by the kernel $K_{s}$ such that $g_{a}(x)\approx f_{s}(x)$ for $x\in\mathcal{M}$?
  • Given a function $g_{a}\in\mathcal{H}_{K_{a}}$ is there a function $f$ in the RKHS $\mathcal{H}_{K}$ defined by the kernel $K$ such that $f(x)\approx 0$ for $x\in\mathcal{M}$ and such that its $f_{a}$ mode is $g_{a}$ and its $f_{z}$ mode is zero?
Then, the natural answer to the questions is to identify the modes of the constraint $f=f_{a}+f_{s}+f_{z}\in\mathcal{H}$ (such that $f(x)\approx 0$ for $x\in\mathcal{M}$ ) such that $f_{a}=-g_{a}$ and $f_{z}=0$ by selecting $f_{s}$ to be the minimizer of the following variational problem
@@
Taking $g_{a}(x)=x_{1}$ for our running example 1, the previous questions are, as illustrated in Fig. 2(b), equivalent to asking whether there exists a function $f_{s}\in\mathcal{H}_{K_{s}}$ that does not depend on $x_{1}$ (since $K_{s}$ does not depend on $x_{1}$) such that
@@
As our running example 1, we can decompose $K_{s}=$(5.33) as the sum of an affine kernel, a quadratic kernel, and a fully nonlinear kernel, i.e., $m=3$, $K_{s,1}=1+\beta_{1}\sum_{i\not=1}x_{i}x_{i}^{\prime}$, $K_{s,2}=\beta_{2}\sum_{i\leq j,i,j\not=1}x_{i}x_{j}x_{i}^{\prime}x_{j}^{\prime}$ and $K_{s,3}=\beta_{3}\prod_{i\not=1}(1+k(x_{i},x_{i}^{\prime}))$.
@@
Then, we can order these sub-modes from most active to least active and create a new kernel $K_{s}$ by removing the least active modes from the signal and adding them to the mode that is set to be zero (see Fig. 13). To describe this, let $\pi(1),\cdots,\pi(m)$ be an ordering of the modes by their activation, i.e., $\|f_{s,\pi(1)}\|_{K_{s,\pi(1)}}^{2}\geq\|f_{s,\pi(2)}\|_{K_{s,\pi(2)}}^{2}\geq\cdots$.
@@
For our running example 1, where we try to identify the ancestors of the variable $x_{1}$, if the sub-mode associated with the variable $x_{2}$ is found to be least active, then we can try to remove $x_{2}$ from the list of ancestors and try to identify $x_{1}$ as a function of $x_{3}$ to $x_{d}$. This is equivalent to selecting $K_{a}(x,x^{\prime})=\beta_{1}x_{1}x_{1}^{\prime}$,
@@
Our initial approach to determining the list of ancestors of a given node is to use a fixed threshold (e.g., $\tau=0.5$) to prune nodes. We propose a refined approach that mimics the strategy employed in Principal Component Analysis (PCA) for deciding which modes should be kept and which ones should be removed. The PCA approach is to order the modes in decreasing order of eigenvalues/variance and (1) either keep the smallest number modes holding/explaining a given fraction (e.g., $90\%$) of the variance in the data, (2) or use an inflection point/sharp drop in the decay of the eigenvalues to select which modes should be kept. Here, we propose a similar strategy. First we employ an alternative determination of the least active mode: we iteratively remove the mode that leads to the smallest increase in noise-to-signal ratio, i.e., we remove the mode $t$ such that,
@@
For our running example 1 in which we try to find the ancestors of the variable $x_{1}$ this is equivalent to removing the variables or node $t$ whose removal leads to the smallest loss in signal-to-noise ratio (or increase in noise-to-signal ratio) by selecting
@@
Next, we iterate this process, and we plot (a) the noise-to-signal ratio, and (b) the increase in noise-to-signal ratio as a function of the number of ancestors ordered according to this iteration. Fig. 14 illustrates this process and shows that the removal of an essential node leads to a sharp spike in increase in the noise-to-signal ratio (the noise-to-signal ratio jumps from approximately 50-60% to 99%). The identification of this inflection point can be used as a method for effectively and reliably pruning ancestors.
Figure 14. Computing the ancestors of the variable $\dot{x}_{0}$ in the Fermi-Pasta-Ulam-Tsingou problem. (a) Noise-to-Signal Ratio, denoted as $\frac{\mathcal{V}(n)}{\mathcal{V}(s)+\mathcal{V}(n)}(q)$, with respect to the number of proposed ancestors, represented by $q$. Additionally, we include a visualization of the quantiles derived from the $Z$-test, as described in Section 7.3. Notably, when there is no signal present, the noise-to-signal ratio is expected to fall within the shaded area with a probability of 0.9. (b) Increments in the Noise-to-Signal Ratio, defined as $\frac{\mathcal{V}(n)}{\mathcal{V}(s)+\mathcal{V}(n)}(q)-\frac{\mathcal{V}(n)}{\mathcal{V}(s)+\mathcal{V}(n)}(q-1)$, as a function of the number of ancestors, denoted as $q$. The horizontal axis represents the number of proposed ancestors for $\dot{x}_{0}$. Determining an appropriate stopping point based solely on absolute noise-to-signal ratio levels can be challenging. In contrast, the increments in the noise-to-signal ratio clearly exhibit a discernible maximum, offering a practical point for decision-making.
@@
2:Learned hypergraph // Set of ancestors for each node
3:$D$ $\leftarrow$ NormalizeData($D$) // Normalize the data
4:for $v\in V$ do
5:  for kernel $\in$ [“linear”, “quadratic”, “nonlinear”] do // Find the kernel
6:   SetOfAncestors($v$) $\leftarrow$ All other nodes
@@
8:   if SignalToNoiseRatio $\tau$ then choose that kernel and exit the for loop
@@
12:  while SignalToNoiseRatio $\tau$ do // Prune ancestors
@@
15:   if SignalToNoiseRatio $\tau$ then Remove that ancestor
@@
Our overall method is summarized in the pseudocode Alg. 1 and Alg. 2 that we will now describe. Alg. 1 takes the data $D$ (encoded into the samples $X_{1},\ldots,X_{N}$ of Problem 1) and the set of nodes $V$ as an input and produces, as described in Sec. 5.3, for each node $i\in V$ its set of minimal ancestors $A_{i}$ and the simplest possible function $f_{i}$ such that $x_{i}\approx f_{i}\big((x_{j})_{j\in A_{i}}\big)$. It employs the default threshold of $0.5$ on the signal-to-noise ratios for its operations. Line 3 normalizes the data (via an affine transformation) so that the samples $X_{i}$ are of mean zero and variance $1$. Given a node with index $i=1$ in Line 4 ($i$ runs through the set of nodes, and we select $i=1$ for ease of presentation), the command in Line 5 refers to selecting a signal kernel of the form $K_{s}=$(5.33) (where $k$ is selected to be a vanilla RBF kernel such as Gaussian or Matérn), with $1\geq\beta_{1}>0=\beta_{2}=\beta_{3}$ for the linear kernel, $1\geq\beta_{1}\geq\beta_{2}>0=\beta_{3}$ for the quadratic kernel and $1\geq\beta_{1}\geq\beta_{2}\geq\beta_{3}>0$ for the fully nonlinear (interpolative) kernel. The ComputeSignalToNoiseRatio function in Line 7 computes the signal-to-noise ratio with $g_{a}(x)=x_{1}$ and with the kernel selected in Line 5. The value of $\gamma$ is selected automatically by maximizing the variance of the histogram of eigenvalues of $D_{\gamma}$ as described in Sec. 7.2 (with the kernel $K=K_{s}=$(5.33) selected in Line 5 and $Y=g_{a}(X)$ with $g_{a}(x)=x_{1}$). The value of $\gamma$ is re-computed whenever a node is removed from the list of ancestors, and $K_{s}$ is nonlinear. Lines 13, 14 and 15 are described in Sec. 5.3.5. They correspond to iteratively identifying the ancestor node $t$ contributing the least to the signal and removing that node from the set of ancestors of the node $1$ if the removal of that node $t$ does not send the signal-to-noise ratio below the default threshold $0.5$.
@@
2:Learned hypergraph // Set of ancestors for each node
3:$D$ $\leftarrow$ NormalizeData($D$) // Normalize the data
4:for node $v\in V$ do
5:  for kernel $\in$ [“linear”, “quadratic”, “nonlinear”] do // Find the kernel
@@
8:   if SignalToNoiseRatio $\tau$ then choose that kernel and exit the for loop
@@
13:  SetOfAncestors(q)(q) $\leftarrow$ All other nodes
@@
17:   SetOfAncestors(q1)(q-1) $\leftarrow$ SetOfAncestors($q$) $\setminus$ LeastImportantAncestor
18:   $q$ $\leftarrow$ q1q-1
@@
20:  $q^{\dagger}$ $\leftarrow$ Inflection point in ($q\rightarrow$ NoiseToSignalRatio(q)(q)) or spike in ($q\rightarrow$ NoiseToSignalRatio(q)(q) - NoiseToSignalRatio(q1)(q-1))
[…]